OpenStudy (anonymous):

find the slope intercept form for the line satisfying the following conditions. x intercept 2, y intercept 2/3

OpenStudy (anonymous):

this gives you two points (2,0) and (0,2/3)

OpenStudy (tkhunny):

The Interecpt Form \(\dfrac{x}{2} + \dfrac{y}{2/3} = 1\) After a little algebra...

OpenStudy (anonymous):

the slope is 1?

OpenStudy (tkhunny):

Who said that? You task is to find the Slope-Intercept Form. @tomo showed you taht you have two points. I showed you the Intercept Form. Now, go! Find the Slope Intercept form.

OpenStudy (anonymous):

@tkhunny and @tomo yeah, I don't understand it either. How is she supposed to do that with a fraction under another fraction (the y/2/3)??? I just would like to know how's that's done.

OpenStudy (tkhunny):

Please do not tell me that you do not know how to multiply of divide fractions. \(\dfrac{x}{2} + \dfrac{y}{2/3} = 1\) Reciprocal and Multiply \(\dfrac{x}{2} + \dfrac{3y}{2} = 1\) Now, can you solve for y?

OpenStudy (tkhunny):

Where did the equal sign go? Why doesn't it say "y = "? \(\dfrac{x}{2} + \dfrac{3y}{2} = 1\) Multiply by 2 \(x + 3y = 2\) Now, can you solve for y?

OpenStudy (anonymous):

@tkhunny Divide both sides by 2 and get 1??? \[ \frac{ 3y }{ 2 } = \frac{ 2 }{ 2 }\] =1 Is that correct?

OpenStudy (anonymous):

so, x = 1???? @tkhunny

OpenStudy (tkhunny):

Where did the x go? Where did the y go? x+3y=2 Subtract x from both sides. 3y = -x + 2 Now can you solve for y? Why are you struggling with this? You should have solved about 400 similar equations for a given variable by now. Is this your first? Are you in the right class? Is your online format providing no informatino whatsoever? Seriously, something is not right, here. I blame your teacher or your curriculum or your format - not you. You should be given the tools you need. It appears you don't have them.

OpenStudy (anonymous):

I just... don't understand. I'm sorry if your ticked @tkhunny . :) I'll stop. Hope the girl that asked this question figure this out! :D @Jasmineplayhouse Have fun!

OpenStudy (tkhunny):

If I have any anger, it is with those who haven't prepared you with the right tools. All I'm saying is that there is a disconnect, here. Let's figure out where it is. You may be in the wrong class. I'm not sure. Have you solved equations before? 3x = 6 Can you solve that for x? What value for x makes hte equality true? 2y - 5 = 3 Can yousolve that for y? What value for y makes the equality true? Let's see where you are?